OpenStudy (anonymous):

Tricky one here. As x tends to infinity, lim(sqrt(x^2+6x+3)-sqrt(x^2+mx+5))=1. Find the value of m.

OpenStudy (anonymous):

m=4

OpenStudy (anonymous):

the limit is 0.

OpenStudy (anonymous):

how did you get m=4?

OpenStudy (anonymous):

Simply Multiply the function by its conjunction (up and Down) , then the x^2 terms will cancel out, after that divide it by x term (up and down,) then apply infinity, voilah, u'll get a simple eqn, solve it m=4

OpenStudy (anonymous):

there is no conjunction. its \[\lim_{x \rightarrow \infty} \sqrt{x^{2} + 6x + 3} - \sqrt{x^{2} +mx + 5} = 1\] and that limit is 0.

OpenStudy (anonymous):

no , there is a conjunction, \[\sqrt{x2+6x+3} + \sqrt{x2+mx+5}\]

OpenStudy (anonymous):

ok thanks.

OpenStudy (anonymous):

oh yeah. sorry. XD

OpenStudy (anonymous):

Cool

OpenStudy (nikvist):